OpenStudy (anonymous):

how do you find the y-intercept? From point (3,-4) (2,0) I found the slope which is 4/-1. I have to put it in slope intercept form. how do I figure out the y-intercept?

OpenStudy (anonymous):

y=mx+b substitute in a point and the slope, then solve for b to find y-intercept

OpenStudy (anonymous):

?

OpenStudy (anonymous):

"plug in a slope and point"?

OpenStudy (anonymous):

I have y=4/-1x + ____

OpenStudy (anonymous):

how do you do that?

OpenStudy (anonymous):

use one of the points that were given either (3,-4) or (2,0)

OpenStudy (anonymous):

graph one of those points and do the slope? do you always do tht to find the y-intercept

OpenStudy (anonymous):

is it 8?

OpenStudy (anonymous):

y=mx+b subsitute in the slope y=-4x+b substitute in a point (3,-4) -4=-4(3)+b solve for b b is the y intercept

OpenStudy (anonymous):

and yes it is equal to 8

OpenStudy (anonymous):

oh ok ...thanks a lot!

OpenStudy (anonymous):

slope is -4, select any of the 2 points say (2,0) write the equation in point slope form y- 0 = -4(x -2) simplify y = -4x + 8 slope- intercept form y- intercept is 8